# Unifying Multiple Foundation Models for Advanced Computational Pathology

### v3

@@
Unifying Multiple Foundation Models for Advanced Computational Pathology
Wenhui Lei1,2‡, Yusheng Tan3‡, Anqi Li3‡, Hanyu Chen4‡, Hengrui Tian1, Ruiying Li5, Zhengqun Jiang1, Fang Yan2, Xiaofan Zhang∗1,2, Shaoting Zhang∗1,2
@@
Washington University in St. Louis, St. Louis, USA
The First Hospital of China Medical University, Shenyang, China
University of Science and Technology Beijing, Beijing, China
@@
*Corresponding author: Xiaofan Zhang (xiaofan.zhang@sjtu.edu.cn), Shaoting Zhang (shaoting.zhang@sjtu.edu.cn)
Foundation models have advanced computational pathology by learning transferable visual representations from large histological datasets, yet recent evaluations reveal substantial variability in their performance across tasks. This inconsistency arises from differences in training data diversity and is further constrained by the reliance of many high-performing models on proprietary datasets that cannot be shared or expanded. Offline distillation offers a partial remedy but depends heavily on the size and heterogeneity of the distillation corpus and requires full retraining to incorporate new models. To address these limitations, we propose Shazam, a task-specific online integration framework that unifies multiple pretrained pathology foundation models within a single flexible inference system. Shazam fuses multi-level representations through adaptive expert weighting and learns task-aligned features via online distillation. Across spatial transcriptomics prediction, survival prognosis, tile classification, and visual question answering, Shazam consistently outperforms strong individual models, highlighting its promise as a scalable approach for harnessing the rapid evolution of pathology foundation models in a unified and adaptable manner.
@@ Introduction
Histopathology remains the cornerstone of cancer diagnosis and treatment planning, providing direct visualization of tissue architecture, cellular morphology, and tumor–microenvironment interactions that guide clinical decision-making. Advances in whole-slide imaging (WSI) have transformed this traditionally microscope-based discipline into a scalable digital workflow, enabling the systematic capture, storage, and analysis of gigapixel-resolution slides. This digital transition has laid the foundation for modern computational pathology (CPath), where artificial intelligence systems can interrogate WSI data to support diagnostic, predictive, and prognostic tasks[van2021deep, shmatko2022artificial]. Building on this momentum, recent pathology foundation models have introduced a new paradigm for CPath. By pretraining on millions of tiles extracted from large cohorts of WSIs across diverse organs and cancer types, models such as Virchow[vorontsov2024foundation], UNI[chen2024towards], and H-Optimus-1[hoptimus1] learn broad and transferable representations that enhance data efficiency and improve downstream performance across a wide spectrum of clinical tasks.
@@
Despite these advances, recent benchmarking studies have revealed substantial performance variability among pathology foundation models across tasks[campanella2025clinical, neidlinger2025benchmarking, ma2025pathbench]. Much of this inconsistency reflects differences in the scale, diversity, and institutional origins of the pretraining data, and no single model consistently performs well across diagnostic, prognostic, molecular, and region-level tasks. Compounding this challenge, many high-performing models rely on large proprietary datasets that cannot be shared due to privacy and institutional constraints[xu2024whole, chen2024towards, hoptimus1]. This lack of data accessibility prevents the research community from steadily expanding the pretraining corpus over time, constraining the growth in data diversity that has supported advances in natural image and language foundation models[deng2009imagenet, schuhmann2022laion, commoncrawl_aws].
@@
To address the challenges posed by inaccessible training data, recent work has explored combining existing foundation models through offline knowledge distillation, exemplified by GPFM[ma2025generalizable]. In this framework, multiple pretrained teacher models first generate features on a curated distillation dataset, and a student model is trained to learn from these signals through expert distillation, complemented by self-distillation that aligns local and global representations. While this strategy provides a practical means of integrating knowledge from several models, it faces several inherent limitations. First, because the distillation process is performed offline, incorporating newly developed foundation models requires retraining the student model from scratch, making rapid integration of emerging advances difficult. Second, the effectiveness of offline distillation depends heavily on the scale and diversity of the available distillation dataset, which in practice is restricted to publicly accessible or newly assembled collections. This constraint limits the range of morphologic and clinical variation that the student model can encounter during distillation, reducing its ability to fully capture the complementary behaviors of diverse teacher models[fang2026knowledge]. Third, offline distillation provides limited flexibility for task-specific adaptation, since the student model learns from teacher outputs generated in a task-agnostic manner and cannot selectively emphasize teacher knowledge that is particularly relevant for a specific downstream prediction target. These constraints highlight the need for an alternative strategy that can directly leverage multiple foundation models in a more flexible and adaptive way, without relying on repeated offline retraining or large-scale distillation corpora.
@@
To address these challenges, we developed Shazam, a unified pathology foundation model that integrates multiple pretrained models within a single and flexible inference framework (Fig.1a). In contrast to offline distillation approaches that require assembling a dedicated distillation dataset and retraining a new model, Shazam performs task-specific online integration. It combines multi-level representations from diverse pathology foundation models through a lightweight Mixture-of-Experts (MoE)[fedus2022switch] mechanism and adapts the fusion process directly to the target task. This design enables Shazam to extract the most informative elements from each model without the need to restart pretraining or construct large-scale distillation corpora whenever new foundation models become available. We evaluate Shazam across a wide range of downstream applications, including spatial transcriptomics prediction, whole-slide survival prognosis, tile-level classification, and visual question answering. These tasks span molecular, regional, and slide-level endpoints and together reflect the breadth of clinical challenges encountered in computational pathology. Across these evaluations, Shazam demonstrates consistent and robust performance. In total, it achieves an average ranking score of 1.17 across 30 benchmarks, substantially outperforming the second-best model, Virchow2, which attains an average of 3.20. The results show that flexible online integration of existing foundation models can effectively consolidate their complementary strengths and provide a practical route toward more adaptable and widely applicable computational pathology systems.
@@
Figure 1: a, Overview of the Shazam framework. Multiple pretrained pathology foundation models are used to extract features from WSIs and tiles. For WSI-level tasks, each model first produces slide representations through a multiple-instance-learning module, whereas tile-level tasks directly use tile embeddings. Shazam performs task-specific online feature integration by fusing low-, middle-, and high-level features and assigning adaptive mixture-of-experts weights. A student model is optimized with an online distillation objective to learn cohesive task-aligned representations without the need for offline distillation or retraining. b, Overall ranking of Shazam and baseline foundation models across 30 downstream benchmarks. A lower rank indicates stronger performance, with a rank of 1 assigned to the top-performing model for each task. Shazam attains the best mean ranking and ranks first on 26 of the 30 tasks. c, Comparison of model performance in spatial transcripomics and survival prediction measured by PCC and C-index.
@@ Results
Shazam is built upon five state-of-the-art pathology foundation models, including UNI 2[chen2024towards], Virchow2[zimmermann2024virchow2], H-optimus-1[hoptimus1], Prov-Gigapath[xu2024whole], and Phikon-v2[filiot2024phikonv2largepublicfeature], and is evaluated across a broad spectrum of downstream applications. Our benchmark covers four major task categories: spatial transcriptomics prediction with eight tasks, whole-slide survival analysis with ten tasks, tile-level classification with eleven tasks, and pathology visual question answering. The overall ranking across all thirty tasks is summarized in Fig. 1b. Shazam achieves the strongest average ranking score of 1.17 and ranks first on 26 of the 30 tasks. By comparison, the next best model, Virchow2, attains an average ranking score of 3.20 and leads only 2 tasks. These findings demonstrate that Shazam substantially outperforms the foundation models on which it is built and exhibits strong generalization across diverse molecular, regional, and slide-level prediction settings.
@@ 1 Shazam enhances spatial transcriptomic prediction
Spatial transcriptomics (ST) provides high-resolution maps of gene expression that reveal the molecular architecture of tissues and the tumor microenvironment, offering insights beyond what morphology alone can capture[staahl2016visualization, baccin2020combined, janesick2023high]. However, the high cost of ST assays and the limited availability of standardized paired H&E–ST datasets have traditionally restricted most computational analyses to narrow tasks and small patient cohorts[meylan2022tertiary, bassiouni2023spatial, parigi2022spatial]. Leveraging pathology foundation models to infer spatial gene expression directly from H&E images offers a scalable alternative, enabling the discovery of treatment-relevant biomarkers across much larger datasets. To assess performance on this challenging task, we compared Shazam against a suite of strong pathology foundation-model baselines for predicting spatial gene expression from H&E slides in the HEST-1K dataset[jaume2024hest], focusing on clinically actionable genes as prediction targets. We considered eight organ-specific cohorts with sufficient paired data, including bowel, liver, breast, prostate, skin, brain, kidney, and bladder, and curated clinically actionable gene sets for each cohort. Additional details are provided in Methods and Supplementary Table1.
@@
Across all eight organ types, Shazam consistently achieved the highest PCCs, with an average improvement of 0.08–0.17 over the strongest individual foundation models (Fig. 2a). The baseline models exhibit clear organ-dependent variability, with Virchow2 performing relatively well in brain and skin cancers, H-optimus-1 showing stronger results in prostate and bladder cancers, and UNI 2 occasionally ranking higher in bowel or kidney cancers (details in Supplementary Tables 2,3). These differences indicate that existing pathology foundation models often reflect the tissue compositions and morphological patterns emphasized during pretraining, which makes it challenging for a single model to maintain uniformly strong performance across distinct organs. In contrast, Shazam ranks first in all eight cancer types, achieving an average ranking of 1.0 and significantly outperforming the next-best models, including UNI 2, Virchow2, and H-optimus-1, whose mean ranking scores are approximately 3.25 (pp=0.004). This consistent advantage across glandular, stromal-rich, and neuroepithelial tissues reflects Shazam’s ability to integrate complementary representations from multiple expert models and to provide stable, cross-organ generalization.
@@
Consistent with these quantitative results, Fig. 2b shows a representative bowel cancer case from the HEST-1K dataset. Shazam accurately captured the heterogeneous spatial pattern of IDH1 expression, including the pronounced up-regulation in dysplastic tumor glands relative to adjacent normal glandular epithelium. This tumor-specific elevation is clearly reflected in the ground truth and was faithfully reproduced only by Shazam, whereas baseline models such as UNI 2, Virchow2, H-optimus-1, GigaPath, and Phikon-v2 either failed to distinguish these compartments or produced the opposite trend. Shazam also recovered the expected low-expression domains within stromal and muscular layers, which other models tended to oversmooth or underrepresent. These results illustrate Shazam’s ability to align tissue morphology with gene-level variation, capturing both tumor-associated up-regulation and stromal or muscular down-regulation patterns that single-model architectures frequently mislocalize.
@@
Collectively, these results demonstrate that Shazam effectively learns a unified morphomolecular representation, translating histological features into gene-level spatial expression patterns with superior accuracy and cross-tissue generalization. Taken together, these findings show that Shazam provides a practical and reliable framework for spatial transcriptomic prediction, offering a flexible foundation for downstream molecular analysis without overreliance on costly assays.
@@
Figure 2: a, Quantitative comparison of prediction accuracy across eight organ types in the HEST-1k dataset, measured by Pearson Correlation Coefficient (PCC). Shazam consistently outperforms existing pathology foundation models (UNI 2, Virchow2, H-optimus-1, Phikon-v2, GigaPath). b, Spatial prediction of IDH1 expression in a colorectal cancer sample. Shazam accurately recapitulates the elevated expression within dysplastic tumor glands and the reduced expression in stromal and muscular layers, aligning closely with the measured spatial transcriptomic profile. Baseline models fail to recover these tissue-compartment differences.
@@ 2 Shazam advances survival prognosis
Accurate survival prognosis is central to cancer management, guiding treatment selection, surveillance strategies, and clinical trial allocation[amin2017eighth]. Yet substantial variability in patient outcomes persists even within the same cancer type, underscoring the limitations of existing clinical and genomic biomarkers. Histopathology, obtained routinely at diagnosis and rich in biological information, remains an underleveraged resource for prognostic modeling. To assess whether Shazam’s unified morphomolecular representations can capture outcome-relevant signals, we conducted a pan-cancer evaluation across ten The Cancer Genome Atlas Program (TCGA)[weinstein2013cancer] cohorts. For each cohort, diagnostic whole-slide images were processed to obtain slide-level representations using attention-based multiple instance learning (details in Methods), and Cox proportional hazards models were trained in a five-fold patient-level cross-validation framework to evaluate the generalizability of survival predictions derived directly from H&E slides.
@@
Across the ten cancer types, Shazam achieved the strongest overall prognostic performance, attaining a mean C-Index of 0.666 (Supplementary Tables 4-6). This surpasses the best-performing individual foundation model, H-optimus-1 (0.642), and exceeds the average performance of Virchow2 (0.631), Prov-GigaPath (0.617), UNI 2 (0.617), and Phikon-v2 (0.594). A paired Wilcoxon signed-rank test confirmed that Shazam’s improvement over H-optimus-1 is statistically significant ($p=0.007$), indicating a consistent advantage rather than cohort-specific variability. Shazam delivered noticeable gains in several cancer types with heterogeneous histology, including colon adenocarcinoma (COAD), where the C-Index increased from 0.646 (H-optimus-1) to 0.710, and lung squamous cell carcinoma (LUSC), where it rose from 0.545 (H-optimus-1) to 0.583. These improvements highlight Shazam’s capacity to integrate complementary strengths from multiple foundation models, yielding a more comprehensive and discriminative representation of disease progression than any individual model.
To further evaluate the clinical relevance of these quantitative gains, we performed Kaplan–Meier analyses using Shazam-derived risk scores. Patients were stratified into high-risk and low-risk groups using a median cutoff. As shown in Fig. 3b, Shazam produced clear and statistically significant separation between survival curves in nine of the ten cancer types (two-sided log-rank test, $p<0.05$ in all except CESC). The pronounced divergence between the high- and low-risk groups is consistent with Shazam’s elevated C-Index values and demonstrates its ability to extract outcome-relevant morphologic features from routine histology. These findings collectively show that Shazam’s unified representations not only improve quantitative survival prediction but also support robust patient stratification across diverse cancer types, underscoring its potential value for downstream clinical decision support[el2025whole].
@@
Figure 3: a, C-Index comparison of Shazam and baseline pathology foundation models across multiple cancer cohorts. Shazam consistently outperforms all baselines except in TCGA-CESC, reflecting its enhanced ability to derive prognostic features from histology. b, Kaplan–Meier curves for a representative cohort, showing significant separation between high-risk and low-risk groups predicted by Shazam (two-sided log-rank test, $p<0.05$), except the TCGA-CESC. The observed survival differences highlight the Shazam’s capacity to recover clinically meaningful prognostic information from whole-slide images.

### v4

# Unified integration of pathology foundation models for scalable histopathology analysis

@@
Unified integration of pathology foundation models for scalable histopathology analysis
Wenhui Lei1‡, Yusheng Tan2‡, Anqi Li2‡, Hengrui Tian1, Ruiying Li3, Zhenqun Jiang1, Fang Yan4, Hanyu Chen∗5, Xiaofan Zhang∗1,6, Shaoting Zhang∗1,7
@@
Washington University in St. Louis, St. Louis, USA
University of Science and Technology Beijing, Beijing, China
@@
Department of Surgical Oncology and General Surgery, Key Laboratory of Precision Diagnosis and Treatment of Gastrointestinal Tumours, Ministry of Education, The First Hospital of China Medical University, Liaoning, China
Shanghai Innovation Institute, Shanghai, China
Sensetime Research, Shanghai, China
@@
*Corresponding author: Hanyu Chen (hychen88@cmu.edu.cn), Xiaofan Zhang (xiaofan.zhang@sjtu.edu.cn), Shaoting Zhang (shaoting.zhang@sjtu.edu.cn)
Foundation models have substantially advanced computational pathology by learning transferable visual representations from large histological datasets, yet their performance varies widely across tasks due to differences in training data composition and reliance on proprietary datasets that cannot be cumulatively expanded. Existing efforts to combine foundation models through offline distillation partially mitigate this issue but require dedicated distillation data and repeated retraining to integrate new models. Here we present Shazam, an online integration model that adaptively combines multiple pretrained pathology foundation models within a unified and scalable representation learning paradigm. Our findings show that fusing multi-level features through adaptive expert weighting and online distillation enables efficient consolidation of complementary model strengths without additional pretraining. Across spatial transcriptomics prediction, survival prognosis, tile-level classification, and visual question answering, Shazam consistently outperforms strong individual models, demonstrating that online model integration provides a practical and extensible strategy for advancing computational pathology.
@@ Introduction
Histopathology remains the cornerstone of cancer diagnosis and treatment planning, providing direct visualization of tissue architecture, cellular morphology, and tumor–microenvironment interactions that guide clinical decision-making. Advances in whole-slide imaging (WSI) have transformed this traditionally microscope-based discipline into a scalable digital workflow, enabling the systematic capture, storage, and analysis of gigapixel-resolution slides. This digital transition has laid the foundation for modern computational pathology (CPath), where artificial intelligence systems can interrogate WSI data to support diagnostic, predictive, and prognostic tasks[43, 39]. Building on this momentum, recent pathology foundation models have introduced a new paradigm for CPath. By pretraining on millions of tiles extracted from large cohorts of WSIs across diverse organs and cancer types, models such as Virchow[45], UNI[12], and H-Optimus-1[9] learn broad and transferable representations that enhance data efficiency and improve downstream performance across a wide spectrum of clinical tasks.
@@
Despite these advances, recent benchmarking studies have revealed substantial performance variability among pathology foundation models across tasks[11, 32, 30]. Much of this inconsistency reflects differences in the scale, diversity, and institutional origins of the pretraining data, and no single model consistently performs well across diagnostic, prognostic, molecular, and region-level tasks. Compounding this challenge, many high-performing models rely on large proprietary datasets that cannot be shared due to privacy and institutional constraints[48, 12, 9]. This lack of data accessibility prevents the research community from steadily expanding the pretraining corpus over time, constraining the growth in data diversity that has supported advances in natural image and language foundation models[13, 38, 1].
@@
One line of recent work has sought to mitigate limited access to large-scale pathology training data by combining existing foundation models through offline knowledge distillation, exemplified by GPFM[29]. In this framework, multiple pretrained teacher models first generate features on a curated distillation dataset, and a student model is trained to learn from these signals through expert distillation, complemented by self-distillation that aligns local and global representations. While this strategy provides a practical means of integrating knowledge from several models, it faces several inherent limitations. First, because the distillation process is performed offline, incorporating newly developed foundation models requires retraining the student model from scratch, making rapid integration of emerging advances difficult. Second, the effectiveness of offline distillation depends heavily on the scale and diversity of the available distillation dataset, which in practice is restricted to publicly accessible or newly assembled collections. This constraint limits the range of morphologic and clinical variation that the student model can encounter during distillation, reducing its ability to fully capture the complementary behaviors of diverse teacher models[15]. Third, offline distillation provides limited flexibility for task-specific adaptation, since the student model learns from teacher outputs generated in a task-agnostic manner and cannot selectively emphasize teacher knowledge that is particularly relevant for a specific downstream prediction target. These constraints highlight the need for an alternative strategy that can directly leverage multiple foundation models in a more flexible and adaptive way, without relying on repeated offline retraining or large-scale distillation corpora.
@@
To directly address these challenges, we developed Shazam, a unified ensemble model that integrates multiple pathology foundation models within a single flexible inference system (Fig. 1a). In contrast to offline distillation approaches that require assembling a dedicated distillation dataset and retraining a new model, Shazam performs task-specific online integration. It combines multi-level representations from diverse pathology foundation models through a lightweight Mixture-of-Experts (MoE)[16] mechanism and adapts the fusion process directly to the target task. This design enables Shazam to extract the most informative elements from each model without the need to restart pretraining or construct large-scale distillation corpora whenever new foundation models become available. We evaluate Shazam across a wide range of downstream applications, including spatial transcriptomics prediction, whole-slide survival prognosis, tile-level classification, and visual question answering. These tasks span molecular, regional, and slide-level endpoints and together reflect the breadth of clinical challenges encountered in computational pathology. Across these evaluations, Shazam demonstrates consistent and robust performance. In total, it achieves an average ranking score of 1.17 across 30 benchmarks, substantially outperforming the second-best model, Virchow2[50], which attains an average of 3.20. By enabling task-adaptive, online integration of heterogeneous pretrained models, Shazam provides a practical mechanism for consolidating complementary representations without requiring additional large-scale pretraining or privileged data access. More broadly, this paradigm suggests a shift in pathology foundation modeling from isolated, monolithic models toward modular and compositional systems, offering a scalable path for continual methodological evolution under realistic clinical and institutional constraints.
@@
Figure 1: a, Overview of the Shazam framework. Multiple pretrained pathology foundation models are used to extract features from WSIs and tiles. For WSI-level tasks, each model first produces slide representations through a multiple-instance-learning module, whereas tile-level tasks directly use tile embeddings. Shazam performs task-specific online feature integration by fusing low-, middle-, and high-level features and assigning adaptive mixture-of-experts weights. A student model is optimized with an online distillation objective to learn cohesive task-aligned representations without the need for offline distillation or retraining. b, Overall ranking of Shazam and baseline foundation models across 30 downstream benchmarks. A lower rank indicates stronger performance, with a rank of 1 assigned to the top-performing model for each task. c, Comparison of model performance in spatial transcriptomics and survival prediction measured by PCC and C-index.
@@ Results
Shazam is built upon five state-of-the-art pathology foundation models, including UNI 2[12], Virchow2[50], H-optimus-1[9], Prov-Gigapath[48], and Phikon-v2[18], and is evaluated across a broad spectrum of downstream applications. Our benchmark covers four major task categories: spatial transcriptomics prediction with eight tasks, whole-slide survival analysis with ten tasks, tile-level classification with eleven tasks, and pathology visual question answering. The overall ranking across all thirty tasks is summarized in Fig. 1b. Shazam achieves the strongest average ranking score of 1.17 and ranks first on 26 of the 30 tasks. By comparison, the next best model, Virchow2, attains an average ranking score of 3.20 and leads only 2 tasks. These findings demonstrate that Shazam substantially outperforms the foundation models on which it is built and exhibits strong generalization across diverse molecular, regional, and slide-level prediction settings.
@@ 1 Shazam enhances spatial transcriptomic prediction
Spatial transcriptomics (ST) provides high-resolution maps of gene expression that reveal the molecular architecture of tissues and the tumor microenvironment, offering insights beyond what morphology alone can capture[40, 5, 21]. However, the high cost of ST assays and the limited availability of standardized paired H&E–ST datasets have traditionally restricted most computational analyses to narrow tasks and small patient cohorts[31, 7, 34]. Leveraging pathology foundation models to infer spatial gene expression directly from H&E images offers a scalable alternative, enabling the discovery of treatment-relevant biomarkers across much larger datasets. To assess performance on this challenging task, we compared Shazam against a suite of strong pathology foundation-model baselines for predicting spatial gene expression from H&E slides in the HEST-1k dataset[22], focusing on clinically actionable genes as prediction targets. We considered eight organ-specific cohorts with sufficient paired data, including bowel, liver, breast, prostate, skin, brain, kidney, and bladder, and curated clinically actionable gene sets for each cohort. Additional details are provided in Methods and Supplementary Table1-2.
@@
Across all eight organ types, Shazam consistently achieved the highest PCCs, with an average improvement of 0.08–0.17 over the strongest individual foundation models (Fig. 1a). The baseline models exhibit clear organ-dependent variability, with Virchow2 performing relatively well in brain and skin cancers, H-optimus-1 showing stronger results in prostate and bladder cancers, and UNI 2 occasionally ranking higher in bowel or kidney cancers (details in Supplementary Tables 3,4). These differences indicate that existing pathology foundation models often reflect the tissue compositions and morphological patterns emphasized during pretraining, which makes it challenging for a single model to maintain uniformly strong performance across distinct organs. In contrast, Shazam ranks first in all eight cancer types, achieving an average ranking of 1.0 and significantly outperforming the next-best models, including UNI 2, Virchow2, and H-optimus-1, whose mean ranking scores are approximately 3.25 (pp=0.004). This consistent advantage across glandular, stromal-rich, and neuroepithelial tissues reflects Shazam’s ability to integrate complementary representations from multiple expert models and to provide stable, cross-organ generalization. In addition to cohort-level performance, Fig. 2b presents a gene-wise comparison of PCC values for spatial expression prediction in the HEST bowel cancer cohort. Shazam achieves consistently higher correlations than all baseline models across the majority of target genes, indicating that its performance gains are not driven by a small subset of genes but instead reflect a stable improvement across diverse molecular targets. Similar gene-wise PCC trends are observed in other cancer types and are provided in Supplementary Figs. 3-7.
@@
Consistent with these quantitative results, Fig. 2c shows a representative bowel cancer case from the HEST-1k dataset. Shazam accurately captured the heterogeneous spatial pattern of IDH1 expression, including the pronounced up-regulation in dysplastic tumor glands relative to adjacent normal glandular epithelium. This tumor-specific elevation is clearly reflected in the ground truth and was faithfully reproduced only by Shazam, whereas baseline models such as UNI 2, Virchow2, H-optimus-1, GigaPath, and Phikon-v2 either failed to distinguish these compartments or produced the opposite trend. Shazam also recovered the expected low-expression domains within stromal and muscular layers, which other models tended to oversmooth or underrepresent. These results illustrate Shazam’s ability to align tissue morphology with gene-level variation, capturing both tumor-associated up-regulation and stromal or muscular down-regulation patterns that single-model architectures frequently mislocalize.
@@
Collectively, these results demonstrate that Shazam effectively learns a unified morphomolecular representation, translating histological features into gene-level spatial expression patterns with superior accuracy and
@@
Figure 2: Evaluation of spatial molecular expression prediction across organs and genes. a, Quantitative comparison of prediction accuracy across eight organ types in the HEST-1k dataset, measured by Pearson Correlation Coefficient (PCC). Shazam consistently outperforms existing pathology foundation models (UNI 2, Virchow2, H-optimus-1, Phikon-v2, GigaPath). Box plots show the median (center line), interquartile range (IQR, box limits represent 25th and 75th percentiles), and whiskers extending to 1.5 × IQR; outliers are shown as individual points. b, Gene-wise PCC for spatial expression prediction in the HEST-Bowel cancer cohort, comparing Shazam with baseline pathology foundation models. c, Spatial prediction of IDH1 expression in a colorectal cancer sample. Shazam accurately recapitulates the elevated expression within dysplastic tumor glands and the reduced expression in stromal and muscular layers, aligning closely with the measured spatial transcriptomic profile. Baseline models fail to recover these tissue-compartment differences.
cross-tissue generalization. Taken together, these findings show that Shazam provides a practical and reliable framework for spatial transcriptomic prediction, offering a flexible foundation for downstream molecular analysis without overreliance on costly assays.
@@ 2 Shazam advances survival prognosis
Accurate survival prognosis is central to cancer management, guiding treatment selection, surveillance strategies, and clinical trial allocation[2]. Yet substantial variability in patient outcomes persists even within the same cancer type, underscoring the limitations of existing clinical and genomic biomarkers. Histopathology, obtained routinely at diagnosis and rich in biological information, remains an underleveraged resource for prognostic modeling. To assess whether Shazam’s unified morphomolecular representations can capture outcome-relevant signals, we conducted a pan-cancer evaluation across ten The Cancer Genome Atlas Program (TCGA)[47] cohorts. For each cohort, diagnostic whole-slide images were processed to obtain slide-level representations using attention-based multiple instance learning (details in Methods), and Cox proportional hazards models were trained in a five-fold patient-level cross-validation framework to evaluate the generalizability of survival predictions derived directly from H&E slides.
@@
Across the ten cancer types, Shazam achieved the strongest overall prognostic performance, attaining a mean C-index of 0.666 (Supplementary Tables 5-7). This surpasses the best-performing individual foundation model, H-optimus-1 (0.642), and exceeds the average performance of Virchow2 (0.631), Prov-GigaPath (0.617), UNI 2 (0.617), and Phikon-v2 (0.594). A paired Wilcoxon signed-rank test confirmed that Shazam’s improvement over H-optimus-1 is statistically significant ($p=0.007$), indicating a consistent advantage rather than cohort-specific variability. Shazam delivered noticeable gains in several cancer types with heterogeneous histology, including colon adenocarcinoma (COAD), where the C-index increased from 0.646 (H-optimus-1) to 0.710, and lung squamous cell carcinoma (LUSC), where it rose from 0.545 (H-optimus-1) to 0.583. These improvements highlight Shazam’s capacity to integrate complementary strengths from multiple foundation models, yielding a more comprehensive and discriminative representation of disease progression than any individual model.
To further evaluate the clinical relevance of these quantitative gains, we performed Kaplan–Meier analyses using Shazam-derived risk scores. Patients were stratified into high-risk and low-risk groups using a median
@@
Figure 3: Evaluation of survival prediction across multiple cancer cohorts. a, C-index comparison of Shazam and baseline pathology foundation models across multiple cancer cohorts. Box plots show the median (center line), interquartile range (IQR, box limits represent 25th and 75th percentiles), and whiskers extending to 1.5 × IQR; outliers are shown as individual points. b, Kaplan–Meier curves for a representative cohort, showing significant separation between high-risk and low-risk groups predicted by Shazam (two-sided log-rank test, $p<0.05$), except the TCGA-CESC.
cutoff. As shown in Fig. 3b, Shazam produced clear and statistically significant separation between survival curves in nine of the ten cancer types (two-sided log-rank test, $p<0.05$ in all except CESC). The pronounced divergence between the high- and low-risk groups is consistent with Shazam’s elevated C-index values and demonstrates its ability to extract outcome-relevant morphologic features from routine histology. These findings collectively show that Shazam’s unified representations not only improve quantitative survival prediction but also support robust patient stratification across diverse cancer types, underscoring its potential value for downstream clinical decision support[14].